# The reliable quantum master equation of the Unruh-DeWitt detector

### v2

@@ III Reliable quantum master equation
The derivation of the QMME indicates that as time $t$ approaches infinity (certainly fulfills $t\gg\left|\alpha^{\prime}-\alpha\right|^{-1}$ and $t\gg\tau_{B}$), the errors stemming from the Markov and RWA approximations tend to vanish. However, this mathematical limit is unsatisfactory because, from a physical perspective, it is vague to refer to a ”sufficiently long” timescale or ”oscillate rapidly”. This motivates us to identify a more precise timescale at which the QMME description of open dynamics becomes physically meaningful and reliable. In this section, we analyze the errors that accumulate over time due to the Markov and RWA approximations. By the proposed relaxed van Hove limit, we establish the criteria for the reliability of the Markov approximation and for the applicability of the RWA, which are upheld in a narrower parameter space for the UDW detector.
@@ III.1 General consideration
In his seminal works sec1-13 ; sec1-14 , Davies aimed to provide a mathematically rigorous definition of the quantum Markovian master equation within a general context. He noted that the order of Eq.(7) is $g^{2}$, indicating that by taking the weak coupling limit $g\rightarrow 0$, the evolution of the open system only becomes drastic once $t$ approaches $\mathcal{O}\left(1/g^{2}\right)$. For $t$ earlier than this timescale, however, the perturbative evolution is moderate under the weak coupling limit, as illustrated in Fig.1. Consequently, Davies proposed a genuine limit (later known as the van Hove limit sec3-1 ): $g^{2}t\sim\mathcal{O}(1)$ while simultaneously taking $g\rightarrow 0$ and $t\rightarrow\infty$, under which Eq.(7) converges to the QMME (23) if the correlation function of the quantum field is integrable [^1] Strictly, this means that $\int_{0}^{\infty}|\mathcal{W}(\tau)|(1+\tau)^{\varepsilon}d\tau<\infty$, for certain $\varepsilon>0$, which is a rigorous manifestation of the previously mentioned requirement for the Wightman function to ”decay sufficiently fast”..
We note that Davies’ result is a limit theorem asserting that for a sufficiently small coupling constant, the master equation provides a good approximation of the real dynamics. However, from a physical perspective, it is unsatisfactory, as Davies’ theorem lacks refined conditions to quantify what is meant by a given physical coupling being ”small enough” plus-alicki or, equivalently, a timescale being ”sufficiently long” whenever the prediction of QMME is reliable.
@@ III.2 Reliability conditions for the Markov approximation
To estimate the error introduced by the Markov approximation, we compare Eq.(11), the dynamical equation under the Born approximation, with the Redfield equation Eq.(13) derived after applying the Markov approximation. The error resulting from the Markov approximation can be isolated and reformulated into:
$\frac{d\rho_{d}(t)}{dt}$ $=-g^{2}\int_{0}^{t}d\tau\left\{\mathcal{W}(\tau)\left[\mathfrak{M}(t),\mathfrak{M}(t-\tau)\rho_{d}(t-\tau)\right]+h.c.\right\}$ (26)
$=\underbrace{-g^{2}\int_{0}^{\infty}d\tau\left\{\mathcal{W}(\tau)\left[\mathfrak{M}(t),\mathfrak{M}(t-\tau)\rho_{d}(t)\right]+h.c.\right\}}_{\text{Markov }}$
$\underbrace{+g^{2}\left\{\int_{0}^{\infty}d\tau\mathcal{W}(\tau)\left[\mathfrak{M}(t),\mathfrak{M}(t-\tau)\left(\rho_{d}(t)-\rho_{d}(t-\tau)\right)\right]+\int_{t}^{\infty}d\tau\mathcal{W}(\tau)\left[\mathfrak{M}(t),\mathfrak{M}(t-\tau)\rho_{d}(t-\tau)\right]\right\}+h.c.}_{\text{error part}}.$
@@
Along the evolution history, we expand $\rho_{d}(t-\tau)$ as:
@@
After some calculations, Eqs. (28) become:
@@
where for brevity, we introduce the notations:
@@
By iterating Eqs. (29) and truncating to order $g^{4}$, we obtain:
@@
To determine the valid range of the Markov approximation, we require that the influence of the first-order error term in Eqs.(28) is sufficiently smaller than the other ”Markov” terms in the equation. Then from Eqs.(31), we obtain following constraints:
@@ III.3 Applicability conditions for the RWA
Under the restrictions (32), Eqs.(31) become:
@@
To demonstrate this, we can recast Eqs.(34) into the GKSL form:
@@
If Eq.(35) is the correct QMME, like (23), the Kossakowski matrix must be Hermitian and positive sec1-13 ; sec1-14 . However, this cannot be true since among the three eigenvalues of the matrix (36):
@@
$\lambda_{3}$ is always negative. This proves that Eqs. (34) cannot preserve the CP property, i.e., the density matrix can become non-positive and, thus, physically unreliable. In other words, the RWA is still necessary to convert Eqs. (34) to a correct QMME [^2] While our analysis aligns with Davies’ theorems, yielding $g$-dependent reliable conditions for Markov and RWA approximations. It is worth noting that series studies sec1-33 ; sec1-34 ; sec1-35 ; sec1-36 offer $g$-independent conditions from an effective field theory (EFT) perspective, where negative $\lambda_{3}$ approaches zero, thereby resolving the non-CP problem without employment RWA..
@@
We are now positioned to compare the solutions of Eqs. (38) and Eqs.(34). First, we note that the second equation of (38) are fairly straightforward to solve, yielding:
@@
Analytically, we can resolve Eq.(41) and reverting it back to the interaction picture gives:
@@
By comparing Eq.(39) and Eq.(43), we aim to determine the conditions under which these two equations are equivalent. First, we observe that the dependency on $\rho_{12}^{*}(0)$ in Eq.(43) disappears in Eq.(39), which implies:
@@
Note that during the derivation, we used the expansion:
@@
By expanding Eq.(45) again using Eq.(47) and neglecting higher-order terms, we ultimately obtain:
@@
Multiplying both sides by $\omega$, these inequalities reveal how the RWA requires constraining $\omega$ by $g^{2}$:
@@
However, when $\omega$ is fixed, the late-time limit also indicates the timescale at which the master equation becomes accurate:
@@
In this sense, even for the detector with a small $\omega$ is (except when $\omega=0$, in which case the interaction is already turned off), as long as the coupling is sufficiently weak or after $t_{\text{RWA}}$, the oscillatory terms characterized by $\omega$ can be effectively recognized as ”rapid oscillatory”.
In conclusion, we have established the reliability conditions of the Markov approximation (32) and the applicability conditions of RWA (49), respectively. These new criteria refine the conventionally adopted descriptions of ”small enough” or ”sufficiently long timescale” (e.g., vaguely $t\gg\left|\alpha^{\prime}-\alpha\right|^{-1}$ and $t\gg\tau_{B}$). The conditions (32) and (49) limit the valid range for the QMME. It is interesting to note that the background quantum field affects the reliability of the QMME through $C_{\Omega}$ and $\Delta_{\Omega}$, which are determined by the field’s Wightman function. This suggests that in applications of QMME to quantum gravity add , the potential compromise of the UV behavior of quantum fields in its reliability requires serious a priori scrutiny.
@@ IV An illustrated example: the Unruh effect of an accelerating detector
We examine the interaction between the detector and a scalar field. The total Hamiltonian takes the same form of Eq. (1):
@@
Here, the detector is modeled as a two-level atom like $H_{d}=\frac{1}{2}\omega\sigma_{3}$, and the free Hamiltonian of the scalar field is given by:
@@
where $m$ is the mass of the field. In terms of annihilation and creation operators $a(k)$ and $a^{\dagger}(k)$, the field operator $\phi(x)$ can be Fourier expanded in the momentum space:
@@
with $\omega_{k}=\left(\mathbf{k}^{2}+m^{2}\right)^{1/2}$. The interaction Hamiltonian $H_{\text{int}}$ describes the coupling between the detector and the field, taking the form $\mathfrak{m}(0)\otimes\phi[x(t)]$, as we have stated in Section II, where $\mathfrak{m}(0)$ is the monopole momentum operator of the detector, generally represented as ($\sigma_{+}+\sigma_{-}$). We establish the initial state of the detector and field as
@@
The Wightman function of the massive field can be calculated as sec1-4-1 ; sec4-0-1 :
@@
known as the Kubo-Martin-Schwinger (KMS) condition sec4-0-2 ; sec4-0-3 . This hallmark feature ensures that the transition rate of the detector from the ground state will ultimately reach equilibrium, from which an effective Unruh temperature proportional to its acceleration can be recognized from the detector’s Planckian spectrum:
@@
To see this, we first solve the quantum Markovian master equation (38):
@@
As time passes, the detector undergoes thermalization due to Unruh radiation. After a sufficiently long duration, by combining Eq.(61) and Eq.(62), we obtain the asymptotic state of the detector, which is a Gibbs state:
@@
It is evident that the asymptotic state relies exclusively on the Unruh temperature or the detector’s acceleration, demonstrating the prediction of the Unruh effect in the ”conventional field approach.”
Now, regarding the accelerating UDW detector model, we return to the theme of this paper: while the open dynamics encodes the thermalization process experienced during the Unruh effect, under what conditions is the dynamics derived from the QMME valid? Utilizing previously established conditions (32) and (48), we can address this issue for both massless and massive backgrounds, respectively.
@@ IV.1 The massless case
To examine the condition (32) and condition (48) for accelerating UDW detector, we first need to know the value of $C_{\Omega}$ and $\Delta_{\Omega}$ which are given as sec1-33 :
@@
It is important to note that $\Delta_{\Omega}$ diverges once the regulator $\varepsilon$ be taken to the rigorous limit as $\varepsilon\rightarrow 0$ in Eq.(65). This indicates that the detector is sensitive to the ultraviolet behavior of the quantum field, which is not surprising since the UDW detector is an effective model and would lose its validity at higher energy scales. Nevertheless, we can admit an alternative interpretation of the regulator $\varepsilon$ as the spatial profile of the detector sec4-1 ; sec4-2 ; sec4-3 ; sec4-4 . A detector with a finite spatial profile would not be sensitive to the short-distance behavior of the quantum field, thus avoiding divergences [^3] However, we must note that such a UV regulator, implemented through the finite detector spatial profile, cannot function at arbitrary scales. In particular, the UV divergences will inevitably reappear when taking the pointlike limit for the detector. . In the following, we use $\varepsilon$ to represent the size of the UDW detector.
@@
We next proceed toward the applicability of the RWA (48) which can be given explicitly by plugging Eq.(64) and Eq.(65) as:
@@
In summary, the open dynamics of an accelerating UDW detector interacting with a massless background can be reliably described by the QMME, with additional constraints on small detector size and low energy gap regions. Furthermore, a sufficiently small coupling constant can effectively suppress the magnitude of the error terms.
@@ IV.2 The massive case
For a massive quantum field, the Wightman function takes an intricate form, as shown in (56). By substituting this into the definition (30), we can straightforwardly give $C_{\Omega}$ as:
@@
The calculation of $\Delta_{\Omega}$ is more subtle. Note that in the limit $t\rightarrow 0$, the massive Wightman function exhibits the same asymptotic form as the massless case:
@@
We can divide the integral in (30) into the divergent and convergent parts of $\mathcal{W}(\tau)$ sec1-33 , respectively:
@@
where the second term in the second step no longer has the singularity, allowing us to safely take $\varepsilon\rightarrow 0^{+}$ and refer to its integral as the ”convergent term”. From Eq. (71), we can see that the value of $\Delta_{\Omega}$ primarily depends on the specific form of $\mathcal{W}_{\text{asymp}}(t)$, which remains unchanged in both the massive and massless cases. Consequently, we have the function $\Delta_{\Omega}$ for massive case:
@@
same as the massless case (65).

### v3

@@ III Reliable quantum master equation
The derivation of the QMME indicates that as time $t$ approaches infinity (certainly fulfills $t\gg\left|\alpha^{\prime}-\alpha\right|^{-1}$ and $t\gg\tau_{B}$), the errors stemming from the Markov and RWA approximations tend to vanish. However, this mathematical limit is unsatisfactory because, from a physical perspective, it is vague to refer to a “sufficiently long” timescale or “oscillate rapidly”. This motivates us to identify a more precise timescale at which the QMME description of open dynamics becomes physically meaningful and reliable. In this section, we analyze the errors that accumulate over time due to the Markov and RWA approximations. By the proposed relaxed van Hove limit, we establish the criteria for the reliability of the Markov approximation and for the applicability of the RWA, which are upheld in a narrower parameter space for the UDW detector.
@@ III.1 General consideration
In his seminal works sec1-13 ; sec1-14 , Davies aimed to provide a mathematically rigorous definition of the quantum Markovian master equation within a general context. He noted that the order of Eq.(7) is $g^{2}$, indicating that by taking the weak coupling limit $g\rightarrow 0$, the evolution of the open system only becomes drastic once $t$ approaches $\mathcal{O}\left(1/g^{2}\right)$. For $t$ earlier than this timescale, however, the perturbative evolution is moderate under the weak coupling limit, as illustrated in Fig.1. Consequently, Davies proposed a genuine limit (later known as the van Hove limit sec3-1 ), $g^{2}t\sim\mathcal{O}(1)$ while simultaneously taking $g\rightarrow 0$ and $t\rightarrow\infty$, under which Eq.(7) converges to the QMME (23) if the correlation function of the quantum field is integrable. [^1] Strictly, this means that $\int_{0}^{\infty}|\mathcal{W}(\tau)|(1+\tau)^{\varepsilon}d\tau<\infty$, for certain $\varepsilon>0$, which is a rigorous manifestation of the previously mentioned requirement for the Wightman function to ”decay sufficiently fast”.
We note that Davies’ result is a limit theorem asserting that for a sufficiently small coupling constant, the master equation provides a good approximation of the real dynamics. However, from a physical perspective, it is unsatisfactory, as Davies’ theorem lacks refined conditions to quantify what is meant by a given physical coupling being “small enough” plus-alicki or, equivalently, a timescale being “sufficiently long” whenever the prediction of QMME is reliable.
@@ III.2 Reliability conditions for the Markov approximation
To estimate the error introduced by the Markov approximation, we compare Eq.(11), the dynamical equation under the Born approximation, with the Redfield equation Eq.(13) derived after applying the Markov approximation. The error resulting from the Markov approximation can be isolated and reformulated into
$\frac{d\rho_{d}(t)}{dt}$ $=-g^{2}\int_{0}^{t}d\tau\left\{\mathcal{W}(\tau)\left[\mathfrak{M}(t),\mathfrak{M}(t-\tau)\rho_{d}(t-\tau)\right]+\text{H. c.}\right\}$ (26)
$=\underbrace{-g^{2}\int_{0}^{\infty}d\tau\left\{\mathcal{W}(\tau)\left[\mathfrak{M}(t),\mathfrak{M}(t-\tau)\rho_{d}(t)\right]+\text{H. c.}\right\}}_{\text{Markov }}$
$\underbrace{+g^{2}\left\{\int_{0}^{\infty}d\tau\mathcal{W}(\tau)\left[\mathfrak{M}(t),\mathfrak{M}(t-\tau)\left(\rho_{d}(t)-\rho_{d}(t-\tau)\right)\right]+\int_{t}^{\infty}d\tau\mathcal{W}(\tau)\left[\mathfrak{M}(t),\mathfrak{M}(t-\tau)\rho_{d}(t-\tau)\right]\right\}+\text{H. c.}}_{\text{error part}}.$
@@
Along the evolution history, we expand $\rho_{d}(t-\tau)$ as
@@
After some calculations, Eqs. (28) become
@@
where for brevity, we introduce the notations,
@@
By iterating Eqs. (29) and truncating to order $g^{4}$, we obtain
@@
To determine the valid range of the Markov approximation, we require that the influence of the first-order error term in Eqs.(28) is sufficiently smaller than the other “Markov” terms in the equation. Then from Eqs.(31), we obtain the following constraints:
@@ III.3 Applicability conditions for the RWA
Under the restrictions (32), Eqs.(31) become
@@
To demonstrate this, we can recast Eqs.(34) into the GKSL form,
@@
If Eq.(35) is the correct QMME, like (23), the Kossakowski matrix must be Hermitian and positive sec1-13 ; sec1-14 . However, this cannot be true since among the three eigenvalues of the matrix (36),
@@
$\lambda_{3}$ is always negative. This proves that Eqs. (34) cannot preserve the CP property, i.e., the density matrix can become nonpositive and, thus, physically unreliable. In other words, the RWA is still necessary to convert Eqs. (34) to a correct QMME. [^2] While our analysis aligns with Davies’ theorems, yielding $g$-dependent reliable conditions for Markov and RWA approximations. It is worth noting that series studies sec1-33 ; sec1-34 ; sec1-35 ; sec1-36 offer $g$-independent conditions from an effective field theory (EFT) perspective, where negative $\lambda_{3}$ approaches zero, thereby resolving the non-CP problem without employment RWA.
@@
We are now positioned to compare the solutions of Eqs. (38) and Eqs.(34). First, we note that the second equation of (38) is fairly straightforward to solve, yielding,
@@
Analytically, we can resolve Eq.(41) and reverting it back to the interaction picture gives
@@
By comparing Eq.(39) and Eq.(43), we aim to determine the conditions under which these two equations are equivalent. First, we observe that the dependency on $\rho_{12}^{*}(0)$ in Eq.(43) disappears in Eq.(39), which implies
@@
Note that during the derivation, we used the expansion,
@@
By expanding Eq.(45) again using Eq.(47) and neglecting higher-order terms, we ultimately obtain,
@@
Multiplying both sides by $\omega$, these inequalities reveal how the RWA requires constraining $\omega$ by $g^{2}$,
@@
However, when $\omega$ is fixed, the late-time limit also indicates the timescale at which the master equation becomes accurate,
@@
In this sense, even for the detector with a small $\omega$ is (except when $\omega=0$, in which case the interaction is already turned off), as long as the coupling is sufficiently weak or after $t_{\text{RWA}}$, the oscillatory terms characterized by $\omega$ can be effectively recognized as “rapid oscillatory”.
In conclusion, we have established the reliability conditions of the Markov approximation (32) and the applicability conditions of RWA (49), respectively. These new criteria refine the conventionally adopted descriptions of “small enough” or “sufficiently long timescale” (e.g., vaguely $t\gg\left|\alpha^{\prime}-\alpha\right|^{-1}$ and $t\gg\tau_{B}$). The conditions (32) and (49) limit the valid range for the QMME. It is interesting to note that the background quantum field affects the reliability of the QMME through $C_{\Omega}$ and $\Delta_{\Omega}$, which are determined by the field’s Wightman function. This suggests that in applications of QMME to quantum gravity add , the potential compromise of the UV behavior of quantum fields in its reliability requires serious a priori scrutiny.
@@ IV An illustrated example: the Unruh effect of an accelerating detector
We examine the interaction between the detector and a scalar field. The total Hamiltonian takes the same form of Eq. (1),
@@
Here, the detector is modeled as a two-level atom like $H_{d}=\frac{1}{2}\omega\sigma_{3}$, and the free Hamiltonian of the scalar field is given by
@@
where $m$ is the mass of the field. In terms of annihilation and creation operators $a(k)$ and $a^{\dagger}(k)$, the field operator $\phi(x)$ can be Fourier expanded in the momentum space,
@@
with $\omega_{k}=\left(\mathbf{k}^{2}+m^{2}\right)^{1/2}$. The interaction Hamiltonian $H_{\text{int}}$ describes the coupling between the detector and the field, taking the form $\mathfrak{m}(0)\otimes\phi[x(t)]$, as we have stated in Sec. II, where $\mathfrak{m}(0)$ is the monopole momentum operator of the detector, generally represented as ($\sigma_{+}+\sigma_{-}$). We establish the initial state of the detector and field as
@@
The Wightman function of the massive field can be calculated as sec1-4-1 ; sec4-0-1
@@
known as the Kubo-Martin-Schwinger (KMS) condition sec4-0-2 ; sec4-0-3 . This hallmark feature ensures that the transition rate of the detector from the ground state will ultimately reach equilibrium, from which an effective Unruh temperature proportional to its acceleration can be recognized from the detector’s Planckian spectrum,
@@
To see this, we first solve the quantum Markovian master equation (38),
@@
As time passes, the detector undergoes thermalization due to Unruh radiation. After a sufficiently long duration, by combining Eq.(61) and Eq.(62), we obtain the asymptotic state of the detector, which is a Gibbs state,
@@
It is evident that the asymptotic state relies exclusively on the Unruh temperature or the detector’s acceleration, demonstrating the prediction of the Unruh effect in the “conventional field approach”.
Now, regarding the accelerating UDW detector model, we return to the theme of this paper; while the open dynamics encodes the thermalization process experienced during the Unruh effect, under what conditions is the dynamics derived from the QMME valid? Utilizing previously established conditions (32) and (48), we can address this issue for both massless and massive backgrounds, respectively.
@@ IV.1 The massless case
To examine the condition (32) and condition (48) for accelerating UDW detector, we first need to know the value of $C_{\Omega}$ and $\Delta_{\Omega}$ which are given as sec1-33
@@
It is important to note that $\Delta_{\Omega}$ diverges once the regulator $\varepsilon$ be taken to the rigorous limit as $\varepsilon\rightarrow 0$ in Eq.(65). This indicates that the detector is sensitive to the ultraviolet behavior of the quantum field, which is not surprising since the UDW detector is an effective model and would lose its validity at higher energy scales. Nevertheless, we can admit an alternative interpretation of the regulator $\varepsilon$ as the spatial profile of the detector sec4-1 ; sec4-2 ; sec4-3 ; sec4-4 . A detector with a finite spatial profile would not be sensitive to the short-distance behavior of the quantum field, thus avoiding divergences. [^3] However, we must note that such a UV regulator, implemented through the finite detector spatial profile, cannot function at arbitrary scales. In particular, the UV divergences will inevitably reappear when taking the pointlike limit for the detector. In the following, we use $\varepsilon$ to represent the size of the UDW detector.
@@
We next proceed toward the applicability of the RWA (48) which can be given explicitly by plugging Eq.(64) and Eq.(65) as
@@
In summary, the open dynamics of an accelerating UDW detector interacting with a massless background can be reliably described by the QMME, with additional constraints on small detector size and low-energy gap regions. Furthermore, a sufficiently small coupling constant can effectively suppress the magnitude of the error terms.
@@ IV.2 The massive case
For a massive quantum field, the Wightman function takes an intricate form, as shown in (56). By substituting this into the definition (30), we can straightforwardly give $C_{\Omega}$ as
@@
The calculation of $\Delta_{\Omega}$ is more subtle. Note that in the limit $t\rightarrow 0$, the massive Wightman function exhibits the same asymptotic form as the massless case,
@@
We can divide the integral in (30) into the divergent and convergent parts of $\mathcal{W}(\tau)$ sec1-33 , respectively,
@@
where the second term in the second step no longer has the singularity, allowing us to safely take $\varepsilon\rightarrow 0^{+}$ and refer to its integral as the “convergent term”. From Eq. (71), we can see that the value of $\Delta_{\Omega}$ primarily depends on the specific form of $\mathcal{W}_{\text{asymp}}(t)$, which remains unchanged in both the massive and massless cases. Consequently, we have the function $\Delta_{\Omega}$ for massive case,
@@
which is the same as the massless case (65).